OpenStudy (anonymous):

Factor completely: 9ab + 12ax − 6b − 8x Prime (3b − 4x)(3a − 2) (3b + 4x)(3a − 2) (3b + 4x)(3a + 2)

OpenStudy (anonymous):

You could just calculate all three of those possibilities and see which one fits.

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

what do you get if you expand C?

OpenStudy (anonymous):

9b+12x)(6a-8

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

What's )(? I think you made a mistake.

OpenStudy (anonymous):

Do you know how to the thing called FOIL to expand bracketed things like this?

OpenStudy (anonymous):

\[(a+b)(c+d)=(a+b)c + (a+b)d=ac + bc + ad + bd\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay. well the thing I wrote above is basically what it is. Do you see how that works? That's what you want to apply to parts a, b, c. What you wrote above, looks quite strange to me. I can't see how you would get that from FOIL.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

while u help me im gonna close this question and ask my other questions and just back to this question

OpenStudy (anonymous):

The thing I wrote above is a result of a property called the Distributive Property. I'll explain why it works, intuitively: This is what the distributive property says you can do, for any numbers a, b, c: \[(a+b)c = ac + bc\] This is because multiplying two numbers is like finding the area of a square:  The whole area of the rectangle (a+b) * c. But you could break up the rectangle:  So then the area is a*c + b*c = (a+b) * c. That's the distributive rule.